OpenStudy (anonymous):

solve for x:cosx-sin3x=cos2x

OpenStudy (anonymous):

Please show me the procedure also

OpenStudy (anonymous):

gonna have to write \[\cos(2x)=1-\sin^2(x)\] and \[\sin(3x)=\sin(x) (3\cos^2(x) - \sin^2(x))\] and then do a raft of algebra